OpenStudy (sarahc):

Solve 3 sqrt x-2 + 7 = 28 x = –89 x = 9 x = 51 x = 79

OpenStudy (solomonzelman):

\[\huge\color{red}{\sqrt[3]{x-2}+7=28}\] OR \[\huge\color{red}{3\sqrt{x-2}+7=28}\]

OpenStudy (sarahc):

what?? which one is the answer.

OpenStudy (solomonzelman):

I am asking which of those (equations I typed in red) is the actual question ?

OpenStudy (sarahc):

The second equation

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 3 \sqrt{x-2}+7=28 }\]\[\huge\color{red}{~~~~~~~~~~~~ -7~~~-7 }\]\[\huge\color{blue}{ 3\sqrt{x-2}=21 }\]\[\huge\color{blue}{ \sqrt{x-2}=7 }\]

OpenStudy (solomonzelman):

square both sides, and add 2 to both sides after that.

OpenStudy (sarahc):

please just give me the answer !! i don't have time

OpenStudy (sarahc):

Thank you!!

OpenStudy (solomonzelman):

You welcome!